OpenStudy (anonymous):

How do I factor x squared plus 3xy plus 2y squared? (I don't need just the answer I need a way to get to the answer.)

OpenStudy (anonymous):

\[=\(x+y)\^\{\2\}\+xy+y \^\{\2\}\=\(x+y)\^\{\2\}\+y(x+y)\=\(x+y)\(x+y+y)\=\(x+y)\(x+\2y)\]

OpenStudy (anonymous):

uh.. the equation thing is not working... but, with the equation: (ax + by)^2 = (ax)^2 + 2abxy + (by)^2 however, you equation is not, what i call "regular" so, you have to find two number that multiplied is 2 and summed is 3 knowing this, the only option is 1 and 2 1x2 = 2 1+2 = 3